# On a generalisation of Cameron’s base size conjecture

### v2

@@ 8. Fixed point ratios for non-parabolic actions
and using Lemma 2.1, we find $|x^{L}|>\frac{1}{2}q^{82}$. Since $x^{L}$ can split into at most 22 $G_{0}$-classes, we get $|x^{G}|>\frac{1}{4}q^{82}$.
If $|H|\leqslant q^{46}$, then the trivial bound ${\rm fpr}(x,G/H)\leqslant|H|/|x^{G}|$ is sufficient. Next, assume that $H$ is of type $E_{7}(q^{1/2})$ and let $H_{0}=G_{0}\cap H$. First suppose $q$ is odd. Inspecting [34, Table 4.1] tells us that $H_{0}=E_{7}(q^{1/2}).2$. Let $\xi:\alpha\mapsto\alpha^{q^{1/2}}$, so $H_{0}=\bar{G}_{\xi}$. If $C_{\bar{G}}(x)$ is connected, then a well-known corollary to the Lang-Steinberg theorem [72, I, 2.7] gives $x^{\bar{G}}\cap\bar{G}_{\xi}=x^{\bar{G}_{\xi}}$ and similarly $x^{\bar{G}}\cap\bar{G}_{\xi^{2}}=x^{\bar{G}_{\xi^{2}}}$. Hence, in view of Proposition 3.5 we have $x^{G_{0}}\cap H=x^{H_{0}}$. On the other hand, if $C_{\bar{G}}(x)$ is disconnected, then the fact that $|C_{\bar{G}}(x)\,:\,C_{\bar{G}}(x)^{0}|=2$, implies that $x^{\bar{G}}\cap H_{0}$ might be a union of two distinct $H_{0}$-classes and the same holds for $G_{0}$. We therefore get
@@
We can now inspect Lübeck’s online data on semisimple classes [65] to inspect which elements have disconnected centraliser in $\bar{G}$ and compute ${\rm fpr}(x,G/H)$ using (8). For example, suppose that $C_{\bar{G}}(x)$ is of type $E_{6}T_{1}$. Then [65] tells us that $C_{\bar{G}}(x)$ is disconnected, and in particular we find that
@@
$|x^{G_{0}}|>\frac{|E_{7}(q)|}{2|E_{6}(q)|(q-1)}=\frac{1}{2}q^{54}$
so the claim holds. The other cases are all similar, so we leave the details to the reader. If $q$ is even, then $G_{0}=\bar{G}_{\sigma}$ and again inspecting [34, Table 4.1] we find that $H_{0}=E_{7}(q^{1/2})$. Using [72, I, 2.7] once again, we deduce that $x^{G_{0}}\cap H_{0}$, so the bounds derived in the case where $q$ is odd are again sufficient.
Finally, if $H=N_{G}(\bar{M}_{\sigma})$ for some $\sigma$-stable subgroup $\bar{M}$ of $\bar{G}$, then one can check that the bounds in Table 14 hold by inspecting the proofs of [25, 4.9, 4.10] and [19, 3.4].
Now suppose that $q=2$. In this case the prime order $\bar{G}_{\sigma}$-classes and the possible centraliser types are recorded in [8, Table 2]. If ${\rm dim}\,x^{\bar{G}}\leqslant 66$, then the claim follows from [51, Theorem 2], so we may assume that ${\rm dim}\,x^{\bar{G}}>66$.
If $|H|\leqslant 2^{46}$, then one can can obtain much stronger bounds than the ones in Table 14 by arguing as in the case $q\geqslant 3$, so we may assume that $H=N_{G}(\bar{M})$ for some $\sigma$-stable subgroup $\bar{M}$ of $\bar{G}$. If $66<{\rm dim}\,x^{\bar{G}}\leqslant 84$, then we can check that ${\rm fpr}(x,G/H)<2^{-19}$ for all the remaining possibilities for $H$ by inspecting the proofs of  [25, 4.9, 4.10] and [19, 3.4]. Now if ${\rm dim}\,x^{\bar{G}}>84$, then by inspecting the same proofs, we again obtain sufficient fixed point ratio bounds that this time depend on the value of $z={\rm dim}\,Z(\bar{D}^{0})$, where $\bar{D}=C_{\bar{G}}(x)$. In particular, we find that ${\rm fpr}(x,G/H)<2^{-21}$ if $z\leqslant 3$ and ${\rm fpr}(x,G/H)<2^{-24}$ if $z>3$. The authors derive those bounds depending on $z$ by appealing to [51, 4.5], which implies
@@
where $\delta(x)={\rm dim}\,x^{\bar{G}}-{\rm dim}(x^{\bar{G}}\cap\bar{M})$ and $l$ is the semisimple rank of $\bar{M}$.
Case 2. $x$ is unipotent: We first consider the case $p=2$. If ${\rm dim}\,x^{\bar{G}}\leqslant 54$, then the claim follows directly by [51, Theorem 2], so we may assume that ${\rm dim}\,x^{\bar{G}}>54$. In this case, $x$ does not lie in the class $A_{1}$ or $A_{1}^{2}$, and so inspecting [59, Table 22.2.2] we find that $|x^{G}|>q^{64}$.
@@
Finally, assume that $H=N_{G}(\bar{M}_{\sigma})$ for some $\sigma$-stable subgroup $\bar{M}$ of $\bar{G}$. If $\bar{M}=A_{1}D_{6}$, then the fusion of $H$-classes in $G$ is described in [25, Table 3] and from this we can compute precise fixed point ratios and verify that the claim holds. For example, suppose that $x$ is in class $A_{1}^{3^{\prime\prime}}$, and write $x=uy$, where $u\in A_{1}$ and $y\in D_{6}$. Then  [25, Table 3] tells us that either $u=1$ and $y$ embeds in $D_{6}$ as an $a^{\prime}_{6}$ involution (see [22, Section 3.5.4] for more information on unipotent involutions in orthogonal groups), or $u\neq 1$ (so $u$ has Jordan decomposition $J_{2}]$ on $\mathbb{F}_{q}^{2}$), and $y$ embeds in $D_{6}$ as a $c_{2}$ involution. We therefore get
@@
Finally, one can check that ${\rm fpr}(x,G/H)<q^{-18}$ for all other possibilities of $\bar{M}$ by inspecting the proofs of [25, 4.9, 4.10] and [19, 3.4].
Now assume that $p>2$. If ${\rm dim}\,x^{\bar{G}}\leqslant 66$, then the bound once again follows from [51, Theorem 2]. Now suppose that ${\rm dim}\,x^{\bar{G}}>66$. If $H$ is of type $A_{1}D_{6}$, then we verify the claim by inspecting [25, 4.10], and in every other case the claim is proved by arguing exactly as in the case $p=2$.
Case 3. $x$ is a field automorphism: If $r=2$, then [51, Theorem 2] gives ${\rm fpr}(x,G/H)\leqslant q^{-22}$, so now assume that $r$ is odd. Here Lemma 2.1 gives
$|x^{G}|=|E_{7}(q)\,:\,E_{7}(q^{1/r})|>\frac{1}{2}q^{-133(1-1/r)}\geqslant\frac{1}{2}q^{-88}.$
If $|H|\leqslant q^{46}$, then we get ${\rm fpr}(x,G/H)<\frac{1}{2}q^{-42}$, so the claim holds. If $H$ is of type $E_{7}(q^{1/2})$, then $x$ also acts as a field automorphism on $E_{7}(q^{1/2})$, and so
@@
In all other cases, one can verify that the bounds in Table 14 are satisfied by inspecting the proofs of [25, 4.9, 4.10] and [19, 3.4], so the proof is complete. ∎
@@
First note that either $|H|\leqslant q^{22}$, or $H$ is one of the possibilities recorded in [25, 4.22]. The case $q=2$ will be dealt with using different methods, so for now we will assume $q\geqslant 3$.
Case 1. $x$ is semsimple: If $C_{\bar{G}_{\sigma}}(x)=B_{4}$, then the bound immediately follows from [51, Theorem 2], so it is left to deal with the remaining possibilities for $x$. First assume that $|H|\leqslant q^{22}$. If $C_{\bar{G}_{\sigma}}(x)$ has a $B_{3}$ or $C_{3}$ factor, then ${\rm dim}\,x^{\bar{G}}\geqslant 28$, and inspecting [65] gives
@@
If $H$ is of type $F_{4}(q^{1/2})$, then we refer to [33, Table 8] and we find that $H_{0}=H\cap G_{0}=F_{4}(q^{1/2})$. Now we note that $\bar{G}$ is simply connected, and so $C_{\bar{G}}(x)$ is connected. Therefore, appealing to [72, I, 2.7] we find that $x^{G}\cap H_{0}=x^{H_{0}}$. We can therefore compute precise fixed point ratios for all semisimple elements and verify that the bounds in Table 15 are satisfied for this choice of $H$. In all other cases we appeal to the proofs of [25, 4.24–4.27] and we verify that the claim holds.
Case 2. $x$ is unipotent, field or graph-field automorphism: The bounds in Table 15 can be directly verified by inspecting [25, 4.23–4.27], unless $H$ is of type $F_{4}(q^{1/2})$. If $x$ is unipotent, then we claim that $x^{G}\cap H=x^{H_{0}}$. This follows from the fact that both the $G_{0}$ and $H_{0}$-class of $x$ are determined by the labelling of its class in $\bar{G}$, and so one can verify that the bounds in Table 15 are satisfied. Finally, if $x$ is a field (respectively, graph-field) automorphism, then $x$ also acts as a field (respectively, graph-field) automorphism on $H$, and so we get
@@
Now assume that $q=2$, so $G=G_{0}$, or $G_{0}.2$. If $x$ is an involutory graph-field automorphism, then [51, Theorem 2] implies that ${\rm fpr}(x,G/H)<2^{-6}$, so we may now assume that $x\in G_{0}$.
Recall that if $|H|>2^{22}$, then the possibilities for $H$ are determined in [25, 4.22], and so we can check that the character tables of both $G_{0}=F_{4}(2)$ and $H_{0}=H\cap G_{0}$ are computable in GAP. It follows that we can determine the fusion of $H_{0}$-classes in $G_{0}$ and we note that ${\rm fpr}(x,G/H)\leqslant\frac{|x^{G_{0}}\cap H_{0}|}{|x^{G_{0}}|}$ by Lemma 2.4. Now assume that $|H|\leqslant 2^{22}$. If ${\rm dim}\,x^{\bar{G}}<28$, then as noted in [25, 4.23] $x$ is an involution belonging to one of the $\bar{G}$-classes labelled $A_{1},\tilde{A}_{1}$, or $\tilde{A}_{1}^{(2)}$, and we have ${\rm fpr}(x,G/H)\leqslant(2^{4}-2^{2}+1)^{-1}=13^{-1}$ by [51, Theorem 1] if $x\in A_{1}$ or $\tilde{A}_{1}$, and ${\rm fpr}(x,G/H)<2^{-4}$ if $x\in\tilde{A}_{1}^{(2)}$. If ${\rm dim}\,x^{\bar{G}}\geqslant 28$, then ${\rm fpr}(x,G/H)\leqslant 2^{22}/|x^{G}|$ We now set
@@ 9. Fixed point ratios for parabolic actions
In this section, we describe how we obtain sufficient fixed point ratio estimates for groups with socle in $G_{0}$ over a given parabolic subgroup $H$ of $G$ when $G_{0}=\mathcal{E}_{2}\cup\mathcal{E}_{3}$. We will be using the same methods that are used and discussed in [51]. In particular, if $\Omega=G/H$, then we have
@@
where $\chi(x)=1_{\bar{H}_{\sigma}}^{\bar{G}_{\sigma}}$ is the corresponding permutation character. We divide the section into four subsections, where we discuss our approach for semisimple elements, unipotent elements, field and graph-field automorphisms, and graph automorphisms respectively.
@@ 9.1. Semisimple elements
Let $W$ denote the Weyl group of $\bar{G}$ and let $W_{\bar{H}}$ denote the Weyl group of $\bar{H}$, so $W_{\bar{H}}$ is a standard parabolic subgroup of $W$. We write $\Phi$ for the root system of $\bar{G}$ with respect to a fixed maximal torus, we let $\Pi$ be a simple system of roots for $\bar{G}$ and $\alpha_{0}$ to be the highest root in $\Phi$ with respect to $\Pi$. The possible centraliser types of semisimple elements in $\bar{G}_{\sigma}$ are in one-to-one correspondence with pairs $(J,[w])$, where $J\subsetneq\Pi\cup\{\alpha_{0}\}$ and is determined up to conjugacy in $W$ and $w]=W_{J}w$ is a conjugacy class representative of $N_{W}(W_{J})/W_{J}$, where $W_{J}$ denotes the subgroup of $W$ generated by reflections in the roots in $J$. This is all described in [36, 39, 40].
Recall that $W$ is $\sigma$-stable, and so $\sigma$ induces permutation of the elements of $W$. We say that two elements $w,w^{\prime}\in W$ are $\sigma$-conjugate if there exists $x\in W$ such that $x^{-1}w\sigma(x)$. Now let $C_{1},\ldots,C_{l}$ be the $\sigma$-conjugacy classes of $W$. As explained in [51, Section 3], the $\sigma$-stable maximal tori of $\bar{G}$ are parametrised by the $\sigma$-classes of $W$, so we will write $T_{i}$ for a representative of the $\bar{G}_{\sigma}$-class of maximal tori in $\bar{G}$ corresponding to $C_{i}$ for every $i\in\{1,\ldots,l\}$, so that $(T_{1})_{\sigma},\ldots,(T_{l})_{\sigma}$ is a list of representatives of the $\bar{G}_{\sigma}$-classes of maximal tori of $\bar{G}_{\sigma}$.
@@
Let $x$ be in the $\bar{G}_{\sigma}$-class corresponding to the pair $(J,[w])$. Using the above description of $\bar{G}_{\sigma}$-classes, Lawther, Liebeck, and Seitz provide an explicit formula for the number of fixed points of $x$ in $\Omega$ in [51, 3.2], which we record below:
@@
We finish the discussion on semisimple elements by giving a description of how one can compute fixed point ratios of semisimple elements in Magma using Proposition 9.1 and Corollary 9.2. Suppose that we want to compute the fixed points for a semisimple element $x\in\bar{G}_{\sigma}$ corresponding to the pair $(J,[w])$. The key ingredient is the function fprSemisimple that takes as input a group of Lie type $G$, a set encoding a parabolic subgroup $P$ of $G$ in Magma (which in our case will be $H$), a set $J$, where $J$ is as defined above, a polynomial poly, which is the $p^{\prime}$-part of the centraliser of the element in question, and a word wrd in standard generators for $W$ that encodes $w$ in Magma. Below we provide the relevant version of fprSemisimple for the case where $\bar{G}_{\sigma}$ is untwisted. If $\bar{G}_{\sigma}$ is twisted, then one can obtain a similar function with very minor modifications. Similarly, if $G_{0}=E_{8}(q)$ and we want to use the approximation of $|C_{\Omega}(x)|$ given by Corollary 9.2, it is possible to do so only with slight modifications of the function provided.
@@
Let $x\in G$ be a semisimple element whose $\bar{G}_{\sigma}$-class corresponds to the pair $(J,[w])$ for $J=\{\alpha_{0},\alpha_{1},\alpha_{2}\}$ and $w=a_{1}a_{3}$, as given in [39, p.123]. Translating the Fleischmann and Janiszczak to Bourbaki notation, we find that $a_{1}$ and $a_{3}$ are the reflections corresponding to $\alpha_{7}$ and $\alpha_{5}$ in $W$ respectively. We also see that $C_{\bar{G}_{\sigma}}(x)$ is of type $A_{2}A_{1}T_{4}$, and in particular $|C_{\bar{G}_{\sigma}}(x)|=q^{4}(q^{2}-1)^{4}(q^{3}-1)$. Finally, noting that the $P_{1}$ parabolic of $G$ is given by StandardParabolicSubgroup(W, {2, 3, 4, 5, 6, 7}) in Magma, we can run the following lines of code:

### v3

@@ 8. Fixed point ratios for non-parabolic actions
and using Lemma 2.5, we find $|x^{L}|>\frac{1}{2}q^{82}$. Since $x^{L}$ is a union of at most $2$ distinct $G_{0}$-classes, we get $|x^{G}|>\frac{1}{4}q^{82}$.
If $|H|\leqslant q^{46}$, then the trivial bound ${\rm fpr}(x,G/H)\leqslant|H|/|x^{G}|$ is sufficient. Next, assume that $H$ is of type $E_{7}(q^{1/2})$ and let $H_{0}=G_{0}\cap H$.
First suppose $q$ is odd. Inspecting [35, Table 4.1] tells us that $H_{0}={\rm Inndiag}(E_{7}(q^{1/2}))=E_{7}(q^{1/2}).2$. Let $\xi:\alpha\mapsto\alpha^{q^{1/2}}$, so $H_{0}=\bar{G}_{\xi}$. If $C_{\bar{G}}(x)$ is connected, then a well-known corollary to the Lang-Steinberg theorem [75, I, 2.7] gives $x^{\bar{G}}\cap\bar{G}_{\xi}=x^{\bar{G}_{\xi}}$ and similarly $x^{\bar{G}}\cap\bar{G}_{\xi^{2}}=x^{\bar{G}_{\xi^{2}}}$. Hence, in view of Proposition 3.5 we have $x^{G_{0}}\cap H=x^{H_{0}}$. On the other hand, if $C_{\bar{G}}(x)$ is disconnected, then the fact that $|C_{\bar{G}}(x)\,:\,C_{\bar{G}}(x)^{0}|=2$ implies that $x^{\bar{G}}\cap H_{0}$ might be a union of two distinct $H_{0}$-classes. Similarly, $x^{\bar{G}}\cap G_{0}$ might be a union of two distinct $G_{0}$-classes. We therefore get
@@
We can now inspect Lübeck’s online data on semisimple classes [66] to deduce which elements have disconnected centraliser in $\bar{G}$ and compute ${\rm fpr}(x,G/H)$ using (10).
For example, suppose that $C_{\bar{G}}(x)$ is of type $E_{6}T_{1}$. Then [66] tells us that $C_{\bar{G}}(x)=E_{6}T_{1}.2$ is disconnected, and in particular we find that
@@
$|x^{G_{0}}|=\frac{|E_{7}(q)|}{2|E_{6}(q)|(q-1)}>\frac{1}{2}q^{54}$
so the claim holds. The other cases are all similar, so we leave the details to the reader.
If $q$ is even, then $G_{0}=\bar{G}_{\sigma}$ and again inspecting [35, Table 4.1] we find that $H_{0}=E_{7}(q^{1/2})$. Using [75, I, 2.7] once again, we deduce that $x^{G_{0}}\cap H_{0}=x^{H_{0}}$, so the bounds derived in the case where $q$ is odd are again sufficient.
Finally, if $H=N_{G}(\bar{M}_{\sigma})$ for some $\sigma$-stable subgroup $\bar{M}$ of $\bar{G}$, then one can check that the bounds in Table 15 hold by inspecting the proofs of [26, 4.9, 4.10] and [19, 3.4].
Now suppose that $q=2$. In this case the prime order $\bar{G}_{\sigma}$-classes and the possible centraliser types are conveniently recorded in [8, Table 2]. If ${\rm dim}\,x^{\bar{G}}\leqslant 66$, then the claim follows from [52, Theorem 2], so we may assume that ${\rm dim}\,x^{\bar{G}}>66$.
If $|H|\leqslant 2^{46}$, then one can can obtain much stronger bounds than the ones in Table 15 by arguing as in the case $q\geqslant 3$, so we may assume that $H=N_{G}(\bar{M}_{\sigma})$ for some $\sigma$-stable subgroup $\bar{M}$ of $\bar{G}$. If $66<{\rm dim}\,x^{\bar{G}}\leqslant 84$, then we can check that ${\rm fpr}(x,G/H)<2^{-19}$ for all the remaining possibilities for $H$ by inspecting the proofs of  [26, 4.9, 4.10] and [19, 3.4]. Now assume ${\rm dim}\,x^{\bar{G}}>84$. By inspecting the same proofs, we again obtain sufficient fixed point ratio bounds that this time depend on the value of $z={\rm dim}\,Z(\bar{D}^{0})$, where $\bar{D}=C_{\bar{G}}(x)$. In particular, we find that ${\rm fpr}(x,G/H)<2^{-21}$ if $z\leqslant 3$ and ${\rm fpr}(x,G/H)<2^{-24}$ if $z>3$. The authors derive those bounds depending on $z$ by appealing to [52, 4.5], which implies that
@@
where $W(\bar{G})$ and $W(\bar{M})$ denote the Weyl groups of $\bar{G}$ and $\bar{W}$ respectively, $\delta(x)={\rm dim}\,x^{\bar{G}}-{\rm dim}(x^{\bar{G}}\cap\bar{M})$ and $l$ is the semisimple rank of $\bar{M}$.
Case 2. $x$ is unipotent: We first consider the case $p=2$. If ${\rm dim}\,x^{\bar{G}}\leqslant 54$, then the claim follows directly by [52, Theorem 2], so we may assume that ${\rm dim}\,x^{\bar{G}}>54$. In this case, $x$ does not lie in the class $A_{1}$ or $A_{1}^{2}$, and so inspecting [60, Table 22.2.2] we find that $|x^{G}|>q^{64}$.
@@
Finally, assume that $H=N_{G}(\bar{M}_{\sigma})$ for some $\sigma$-stable subgroup $\bar{M}$ of $\bar{G}$. If $\bar{M}=A_{1}D_{6}$, then the fusion of $H$-classes in $G$ is described in [26, Table 3] and from this we can compute precise fixed point ratios and verify that the claim holds. For example, suppose that $x$ is in class $(A_{1}^{3})^{(2)}$, and write $x=uy$, where $u\in A_{1}$ and $y\in D_{6}$. Then  [26, Table 3] tells us that either $u=1$ and $y$ embeds in $D_{6}$ as an $a^{\prime}_{6}$ involution (see [22, Section 3.5.4] for more information on unipotent involutions in orthogonal groups), or $u\neq 1$ (so $u$ has Jordan decomposition $J_{2}]$ on the natural $A_{1}$-module), and $y$ embeds in $D_{6}$ as a $c_{2}$ involution. We therefore get
@@
Finally, one can check that ${\rm fpr}(x,G/H)<q^{-18}$ for all other possibilities for $\bar{M}$ by inspecting the proofs of [26, 4.9, 4.10] and [19, 3.4].
Now assume that $p>2$. If ${\rm dim}\,x^{\bar{G}}\leqslant 66$, then the desired bound once again follows from [52, Theorem 2]. Now suppose that ${\rm dim}\,x^{\bar{G}}>66$. If $H$ is of type $A_{1}D_{6}$, then we verify the claim by inspecting the proof of [26, 4.10], and in every other case the claim is proved by arguing exactly as in the case $p=2$.
Case 3. $x$ is a field automorphism: If $r=2$, then [52, Theorem 2] gives ${\rm fpr}(x,G/H)\leqslant q^{-22}$, so now assume that $r$ is odd. In this case, we can infer from [35, Table 4.1] that $C_{G_{0}}(x)=E_{7}(q^{1/r})$. Then Lemma 2.5 gives
$|x^{G_{0}}|=|E_{7}(q)\,:\,E_{7}(q^{1/r})|>\frac{1}{2}q^{-133(1-1/r)}\geqslant\frac{1}{2}q^{-88}.$
If $|H|\leqslant q^{46}$, then we get ${\rm fpr}(x,G/H)<\frac{1}{2}q^{-42}$, so the claim holds. If $H$ is of type $E_{7}(q^{1/2})$, then $x$ also acts as a field automorphism on $E_{7}(q^{1/2})$, and so Lemma 2.3 implies
@@
In all other cases, one can verify that the bounds in Table 15 are satisfied by inspecting the proofs of [26, 4.9, 4.10] and [19, 3.4], so the proof is complete. ∎
@@
First note that either $|H|\leqslant q^{22}$, or $H$ is one of the possibilities recorded in [26, 4.23]. The case $q=2$ will be dealt with using different methods, so for now we will assume $q\geqslant 3$.
Case 1. $x$ is semsimple: If $C_{\bar{G}_{\sigma}}(x)=B_{4}$, then the bound immediately follows from [52, Theorem 2], so it is left to deal with the remaining possibilities for $x$. First assume that $|H|\leqslant q^{22}$. If $C_{\bar{G}_{\sigma}}(x)$ has a $B_{3}$ or $C_{3}$ factor, then ${\rm dim}\,x^{\bar{G}}\geqslant 28$, and inspecting [66] gives
@@
If $H$ is of type $F_{4}(q^{1/2})$, then we refer to [34, Table 8] and we find that $H_{0}=H\cap G_{0}=F_{4}(q^{1/2})$. Now we note that $\bar{G}$ is simply connected, and so $C_{\bar{G}}(x)$ is connected. Therefore, appealing to [75, I, 2.7] we find that $x^{G}\cap H_{0}=x^{H_{0}}$. We can therefore compute precise fixed point ratios for all semisimple elements and verify that the bounds in Table 16 are satisfied for this choice of $H$. In all other cases we appeal to the proofs of [26, 4.22, 4.24–4.28] and we verify that the claim holds.
Case 2. $x$ is unipotent, field or graph-field automorphism: The bounds in Table 16 can be directly verified by inspecting [26, 4.22, 4.24–4.28], unless $H$ is of type $F_{4}(q^{1/2})$. If $x$ is unipotent, then we claim that $x^{G_{0}}\cap H=x^{H_{0}}$. This follows from the fact that both the $G_{0}$ and $H_{0}$-class of $x$ are determined by the labelling of its class in $\bar{G}$, and so one can verify that the bounds in Table 16 are satisfied. Finally, if $x$ is a field (respectively, graph-field) automorphism, then $x$ also acts as a field (respectively, graph-field) automorphism on $H$, and so we get
@@
Now assume that $q=2$, so $G=G_{0}$, or $G_{0}.2$. If $x$ is an involutory graph-field automorphism, then [52, Theorem 2] implies that ${\rm fpr}(x,G/H)<2^{-6}$, so we may now assume that $x\in G_{0}$.
Recall that if $|H|>2^{22}$, then the possibilities for $H$ are determined in [26, 4.23], and so we can check that the character tables of both $G_{0}=F_{4}(2)$ and $H_{0}=H\cap G_{0}$ are computable in GAP. It follows that we can determine the fusion of $H_{0}$-classes in $G_{0}$, and we note that ${\rm fpr}(x,G/H)\leqslant\frac{|x^{G_{0}}\cap H_{0}|}{|x^{G_{0}}|}$ by Lemma 2.3. Now assume that $|H|\leqslant 2^{22}$. If ${\rm dim}\,x^{\bar{G}}<28$, then as noted in [26, 4.22], $x$ is an involution belonging to one of the $\bar{G}$-classes labelled $A_{1},\tilde{A}_{1}$, or $\tilde{A}_{1}^{(2)}$, and we have ${\rm fpr}(x,G/H)\leqslant(2^{4}-2^{2}+1)^{-1}=13^{-1}$ by [52, Theorem 1] if $x\in A_{1}$ or $\tilde{A}_{1}$, and ${\rm fpr}(x,G/H)<2^{-4}$ if $x\in\tilde{A}_{1}^{(2)}$. If ${\rm dim}\,x^{\bar{G}}\geqslant 28$, then ${\rm fpr}(x,G/H)\leqslant 2^{22}/|x^{G}|$. We now set
@@ 9. Fixed point ratios for parabolic actions
In this section, we describe how we obtain fixed point ratio estimates for a group $G$ with socle $G_{0}\in\mathcal{E}_{2}\cup\mathcal{E}_{3}$ in the natural action of $G$ on $G/H$ for some maximal parabolic subgroup $H$ of $G$. We will be using the standard $P_{m}$ parabolic notation for maximal parabolic subgroups of exceptional groups. Moreover, we will denote the parabolic subgroup corresponding to removing nodes $m$ and $n$ from the Dynkin diagram of $G$ by $P_{m,n}$.
We will be using the same methods that are used and discussed in [52]. In particular, if $\Omega=G/H$, then we have
@@
where $\chi(x)=1_{\bar{H}_{\sigma}}^{\bar{G}_{\sigma}}$ is the corresponding permutation character. We divide this section into four subsections, where we discuss our approach for semisimple elements, unipotent elements, field and graph-field automorphisms, and graph automorphisms respectively.
@@ 9.1. Semisimple elements
Let $W$ denote the Weyl group of $\bar{G}$ and let $W_{\bar{H}}$ denote the Weyl group of $\bar{H}$, so $W_{\bar{H}}$ is a standard parabolic subgroup of $W$. We write $\Phi$ for the root system of $\bar{G}$ with respect to a fixed maximal torus, we let $\Pi$ be a simple system of roots for $\bar{G}$ and $\alpha_{0}$ to be the highest root in $\Phi$ with respect to $\Pi$. The possible centraliser types of semisimple elements in $\bar{G}_{\sigma}$ are in one-to-one correspondence with the pairs $(J,[w])$, where $J\subsetneq\Pi\cup\{\alpha_{0}\}$ and is determined up to conjugacy in $W$ and $w]=W_{J}w$ is a conjugacy class representative of $N_{W}(W_{J})/W_{J}$, where $W_{J}$ denotes the subgroup of $W$ generated by reflections in the roots in $J$. This is all described in [37, 40, 41].
Recall that $W$ is $\sigma$-stable, and so $\sigma$ induces permutation of the elements of $W$. We say that two elements $w,w^{\prime}\in W$ are $\sigma$-conjugate if there exists $x\in W$ such that $w^{\prime}=x^{-1}w\sigma(x)$. Now let $C_{1},\ldots,C_{l}$ be the $\sigma$-conjugacy classes of $W$. As explained in [52, Section 3], the $\sigma$-stable maximal tori of $\bar{G}$ are parametrised by the $\sigma$-classes of $W$, so we will write $T_{i}$ for a representative of the $\bar{G}_{\sigma}$-class of maximal tori in $\bar{G}$ corresponding to $C_{i}$ for every $i\in\{1,\ldots,l\}$, so that $(T_{1})_{\sigma},\ldots,(T_{l})_{\sigma}$ is a list of representatives of the $\bar{G}_{\sigma}$-classes of maximal tori of $\bar{G}_{\sigma}$.
@@
Let $x$ be in the $\bar{G}_{\sigma}$-class corresponding to the pair $(J,[w])$. Using the above description of $\bar{G}_{\sigma}$-classes, Lawther, Liebeck, and Seitz provide an explicit formula for the number of fixed points of $x$ in $\Omega$ in [52, 3.2], which we record below:
@@
We finish the discussion on semisimple elements by giving a description of how one can compute fixed point ratios of semisimple elements in Magma using Proposition 9.1 and Corollary 9.2. Suppose that we want to compute the fixed points for a semisimple element $x\in\bar{G}_{\sigma}$ corresponding to the pair $(J,[w])$. The key ingredient is the function fprSemisimple that takes as input a group of Lie type $G$, a set encoding a parabolic subgroup $P$ of $G$ in Magma (which in our case will be $H$), a set $J$, where $J$ is as defined above, a polynomial poly, which is the $p^{\prime}$-part of the centraliser of the element in question, and a word wrd in standard generators for $W$ that encodes $w$ in Magma. Below we provide the relevant version of fprSemisimple for the case where $\bar{G}_{\sigma}$ is untwisted. If $\bar{G}_{\sigma}$ is twisted, then one can obtain a similar function with minor modifications. Similarly, if $G_{0}=E_{8}(q)$ and we want to use the approximation of $|C_{\Omega}(x)|$ given by Corollary 9.2, it is possible to do so only with slight modifications of the function provided.
@@
Let $x\in G$ be a semisimple element whose $\bar{G}_{\sigma}$-class corresponds to the pair $(J,[w])$ for $J=\{\alpha_{0},\alpha_{1},\alpha_{2}\}$ and $w=a_{1}a_{3}$, as given in [40, p.123]. Translating the Fleischmann and Janiszczak to Bourbaki notation, we find that $a_{1}$ and $a_{3}$ are the reflections corresponding to $\alpha_{7}$ and $\alpha_{5}$ in $W$ respectively. We also see that $C_{\bar{G}_{\sigma}}(x)$ is of type $A_{2}A_{1}T_{4}$, and in particular $|C_{\bar{G}_{\sigma}}(x)|=q^{4}(q^{2}-1)^{4}(q^{3}-1)$. Finally, noting that the $P_{1}$ parabolic of $G$ is given by StandardParabolicSubgroup(W, {2, 3, 4, 5, 6, 7}) in Magma, we can run the following lines of code: